Circular No. 5915

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993aj IN ANONYMOUS GALAXY Christian Pollas, Observatoire de la Côte d’Azur, reports his discovery

of an apparent supernova (V = 18.5) located at α = 12^{h}01^{m}47^{s}.38, δ = −0^{o}26′31′′.9 (equinox 1950.0), which is 3′′.6 west and 3′′.9 south of the host spiral galaxy (V ∼ 17). The object was found on two Technical Pan films (limiting mag ∼ 20.5) taken 1993 Dec. 27.23 and 28.24 UT. The object is only 1 mm from the side of the field. Nothing is visible on the POSS prints down to mag 19.5-20. A nearby star of mag ∼ 16 is located at α = 12^{h}01^{m}42^{s}.90, δ = −0^{o}24′53′′.9.

M. Hamuy, Cerro Tololo Interamerican Observatory, reports confirmation by J. Storm (European Southern Observatory) of the presence of a new stellar object 4′′.2 west and 4′′.2 south of the galaxy’s center, as observed in B, V, and I images obtained on 1993 Dec. 29.333 UT with the CTIO 0.9-m telescope.

GX 339–4 B. A. Harmon, W. S. Paciesas, S. N. Zhang, G. J. Fishman, and M. H.

Finger, Marshall Space Flight Center, NASA, report for the Compton Observatory’s Burst and Transient Source Experiment (BATSE) Team: “The black-hole candidate GX 339–4 has again entered its hard (low) emission state, being detected by earth occultation beginning in late December 1993. Its intensity has increased gradually and is currently at ∼ 0.2 Crab (Dec. 30-1994 Jan. 1) in the band 20-100 keV with a hard spectrum. Outbursts of GX 339–4 beginning in late June 1991 (IAUC 5327) and early September 1992 (IAUC 5647), and now in December 1993, indicate an interval between hard x-ray outbursts of ∼ 440 ± 30 days. This is a strong indication of long-term periodicity in this suspected-but-unproven binary system. If the periodicity is associated with the binary orbit, GX 339–4 does not fit conveniently into the canonical model of a low-mass, x-ray binary.”

PERIODIC COMET SCHWASSMANN-WACHMANN 1 Further total magnitude estimates (cf. IAUC 5886): 1993 Dec. 4.85

UT, 13.2 (H. Mikuž, Ljubljana, Slovenia, 0.20-m Baker-Schmidt camera + CCD + V filter); 17.98, 14.3 (Mikuž); 24.51, 15.7 (T. Kojima, YGCO Chiyoda Observatory, Japan, 0.25-m reflector + CCD); 29.56, 15.9 (Kojima); 1994 Jan. 2.61, 13.2 (Kojima; comet almost stellar).

1994 January 3 Daniel W. E. Green